1.given the system express y as a function of t

  1. 3dx/dt+2dy/dt-x+y=t-1
  2. dx/dt+dy/dt-x=t+2

2.what do you understand by ordinary paint of the equation

a0(x)d2y/dx2+a1(x)dy/dx+a2(x)y=0? hence using taylor series expansion find a series solution in powers of x for the equation(2x3-3)d2y/dx2-2xdy/dx+y=0 y(0)=-1

y1(0)=5


1
Expert's answer
2021-07-30T10:41:49-0400

1)

Given the system let us express "y" as a function of "t ."


"i)" "3\dfrac{dx}{dt}+2\dfrac{dy}{dt}-x+y=t-1"


"ii)\dfrac{dx}{dt}+\dfrac{dy}{dt}-x=t+2"


Multiplying equation by "\times 3" , we get -


And the adding "i)\ and \ ii)\ ," we get -


"=\dfrac{-dy}{dt}+2x+y=-2t-7"


It gives values of "x" as -


"x=\dfrac{1}{2}\dfrac{dy}{dt}-\dfrac{1}{2}y-t-\dfrac{7}{2}"


Now , differentiating with respect to t , we get -


"=\dfrac{dx}{dt}=" "\dfrac{1}{2}\dfrac{d^{2}y}{dt^{2}}-\dfrac{1}{2}\dfrac{dy}{dt}-1" .


We get the equation as -


"=\dfrac{1}{2}\dfrac{d^{2}y}{dt^{2}}-\dfrac{1}{2}\dfrac{dy}{dt}-1+\dfrac{dy}{dt}-(\dfrac{1}{2}\dfrac{dy}{dt}-\dfrac{1}{2}y-t-\dfrac{7}{2})=t+2"

Which is equivalent to -


"=\dfrac{1}{2}\dfrac{d^{2}y}{dt^{2}}+\dfrac{1}{2}y=\dfrac{-1}{2}"


And hence the equation becomes -


"=\dfrac{d^{2}y}{dt^{2}}+y=-1"


Auxiliary equation becomes -


"=m^{2}+1=0"


It's root's becomes as -


"m=+i,-i" ,


It's follows that solution is of form -


"=y(t)=C_1cost+C_2sint +y_p," where "y_p=a"


It's follows that -


"\dfrac{d^{2}y_p}{dt^{2}}=0" , hence "a=-1." Therefore the solution is the following -


"y(t)=C_1cost +C_2sint-1"


"2)"

Ordinary point of equation is defined as -


A point a is called ordinary point of differential equation when function "p_{1}(x)\ and \ p_0(x)" are analytical at "x=a" ,


"=a_0(x)\dfrac{d^{2}y}{dx^{2}}+a_1(x)\dfrac{dy}{dx}+a_2(x)y=0"


For , the above equation the ordinary point is called as -


A point "x=x_0" is called ordinary point of differential equation if "a_0(x)\neq0" at "x=x_0" . The point "x=x_0" is called that ordinary point .


"=(2x^{3}-3)\dfrac{d^{2}y}{dx^{2}}-2x\dfrac{dy}{dx}+y=0" "y(0)=-1"

"y^{'}(1)=5"


"Now" , we know that taylor's series expansion is expressed as -


"y(x)=" "y(0)+y^{'}(0)(x-0)+\dfrac{1}{2!}y^{''}(0)(x-0)^{2}+...."


"=y^{''}(x)=\dfrac{2xy^{'}(x)}{2x^{3}-3}-\dfrac{y}{2x^{3}-3}"


Now putting , "x=0," we get -


"y^{''}(0)=0-\dfrac{y(0)}{-3}"


"y^{''}(0)=\dfrac{-1}{3}"


"y(x)=-1+5x+\dfrac{1}{2!}\dfrac{-1}{3}x^{2}+........."


This is taylor's series of given equation .